# A simple mechanism for the enhancement of the inflationary power spectrum

### v1

@@ 4 Gravitational waves and primordial black holes
While many GW signals originate from late-time astrophysical sources, an especially intriguing possibility is a primordial contribution from the early Universe [Caprini:2018mtu, Guzzetti:2016mkm]. Scalar perturbations generated during inflation inevitably source gravitational waves at second order, but in minimal slow-roll models the resulting signal is unobservably small [Domenech:2021ztg, Kohri:2018awv]. However, inflationary scenarios with localized enhancements of the curvature power spectrum, such as those produced by the sequences of turns in multi-field trajectories discussed here, can generate a strong stochastic background upon horizon reentry, potentially accompanied by PBH formation [Young:2014ana, Ozsoy2023Review, Figueroa:2023zhu, Balaji:2023ehk].
@@ 4.2 Primordial Black Holes and Observational Constraints
The present-day PBH dark matter fraction, $f_{\rm PBH}(M)\equiv\Omega_{\rm PBH}(M)/\Omega_{\rm DM}$, is proportional to $\beta(M)$. Consequently, the predicted abundance is also highly sensitive to primordial non-Gaussianities, since PBH formation probes the extreme tail of the probability distribution [Byrnes:2012yx, Young:2013oia, Franciolini:2018vbk, Atal:2018neu].
@@ 4.3 Scalar-induced stochastic gravitational waves
In our scenario, strong scalar-induced gravitational waves are mainly produced when the enhanced scalar perturbations reenter the Hubble horizon. During radiation domination, the induced tensor spectrum can be computed within second-order cosmological perturbation theory [Ananda:2006af, Baumann:2007zm, Domenech:2021ztg]. The oscillation-averaged tensor power spectrum takes the form [Kohri:2018awv]

### v2

@@ 4 Gravitational waves and primordial black holes
While many GW signals originate from late-time astrophysical sources, an especially intriguing possibility is a primordial contribution from the early Universe [Caprini:2018mtu, Guzzetti:2016mkm]. Scalar perturbations generated during inflation inevitably source gravitational waves at second order, but in minimal slow-roll models the resulting signal is unobservably small. However, inflationary scenarios with localized enhancements of the curvature power spectrum, such as those produced by the sequences of turns in multi-field trajectories discussed here, can generate a strong stochastic background upon horizon reentry, potentially accompanied by PBH formation [Young:2014ana, Ozsoy2023Review, Figueroa:2023zhu, Balaji:2022dbi, Balaji:2023ehk].
@@ 4.2 Primordial Black Holes and Observational Constraints
The present-day PBH dark matter fraction, $f_{\rm PBH}(M)\equiv\Omega_{\rm PBH}(M)/\Omega_{\rm DM}$, is proportional to $\beta(M)$. Consequently, the predicted abundance is also highly sensitive to primordial non-Gaussianities, since PBH formation probes the extreme tail of the probability distribution [Byrnes:2012yx, Young:2013oia, Franciolini:2018vbk, Atal:2018neu, DeLuca:2019qsy, Firouzjahi:2023xke].
@@ 4.3 Scalar-induced stochastic gravitational waves
In our scenario, strong scalar-induced gravitational waves are mainly produced when the enhanced scalar perturbations reenter the Hubble horizon. During radiation domination, the induced tensor spectrum can be computed within second-order cosmological perturbation theory [Ananda:2006af, Baumann:2007zm, Espinosa:2018eve, Kohri:2018awv, Domenech:2021ztg]. The oscillation-averaged tensor power spectrum takes the form [Kohri:2018awv]